Hemoglobin is the protein that transports O2 through the blood from the lungs to the rest of the body. In doing so, each molecule of hemoglobin combines with four molecules of O2. If 1.37 g of hemoglobin combines with 2.13 mL of O2 at 37°C and 724 torr, what is the molar mass of hemoglobin?


Answer in scientific notation


1
Expert's answer
2021-03-03T06:29:11-0500

Solution:

Gem  + 4O2   =     GemO4

1,37g 2,13ml

n = PV/RT= 0,95*0,00213/0,0821*(273+37)=0.00007950571

x(Gem)=0.00007950571 /4=0.00001987642 mol

xmol=1,37/ 0.00001987642 =68925.8930934

Answer:Gem=68925.8930934g/mol